# Top-quark pair production as a probe of light top-philic scalars and anomalous Higgs interactions

### v1

@@ 5 Higgs boson $H$: Theoretical framework
In Sec. 5.1 we provide the relevant Lagrangian, the notation used and the main formulas that are analogous to the ones provided in Sec. 2 for the case of the scalar $S$. In Sec. 5.2 we explain in details how we have derived such expressions, recycling the results of Sec. 2 for now the Higgs case, highlighting important differences w.r.t. the scalar $S$ case. In Sec. 5.3 we show how this calculation can be reinterpreted in the SMEFT framework, which further support the consistency of our approach.
@@ 5.1 Lagrangian, notation and relevant formulas
$\mathcal{L}_{H,\,{\rm NP}}\equiv-\overline{\psi}_{t}\left[\frac{y_{t}}{\sqrt{2}}+i\frac{\tilde{y}_{t}}{\sqrt{2}}\gamma_{5}\right]\psi_{t}H\,\equiv-\frac{y_{t}^{\rm SM}}{\sqrt{2}}\leavevmode\nobreak\ \overline{\psi}_{t}\left[\left(\kappa_{t}-1\right)+i\tilde{\kappa}_{t}\gamma_{5}\right]\psi_{t}H\,,$ (47)
@@
Unlike the case of the scalar $S$, where the contribution form the $t\bar{t}S$ has been taken into account, we will not include the contribution from the $t\bar{t}H$ final state in our studies for the sensitivity on $\kappa_{t}$ and $\tilde{\kappa}_{t}$ in Sec. 7 as that corresponds to a different analysis at the LHC.
@@ 5.2 Similarities and differences with the scalar $S$ case
From the purely calculation side the two scenarios are therefore equivalent and via Eqs. (56)–(58) it is possible to obtain $\sigma_{\rm NP}^{H}$ in Eq. (51) from $\sigma_{\rm NP}$ in Eq. (31). However, this cannot be achieved by simply applying them directly as we will explain in the next paragraph. As already said, $\sigma_{\rm NP}^{H}$ corresponds to the renormalised one-loop virtual corrections induced by the diagrams of Fig. 1, with $H$ in the place of $S$, to $\sigma_{\rm LO_{\rm QCD}}$, is the LO cross section from purely QCD interactions (the diagrams in Fig. 2). Renormalisation is again understood in the on-shell scheme, as in the scalar S case.
@@ 5.3 The SMEFT perspective
The consistency of the kappa-framework for such calculation and especially the possibility of recycling the calculation for $S$ of Sec. 2.2 also for the Higgs itself and including not only terms proportional to $(\kappa_{t}-1)^{2}\simeq\Re^{2}(C^{u\Phi}_{tt})$ but also to to $\kappa_{t}-1\simeq\Re(C^{u\Phi}_{tt})$ relies on some subtleties that are typically ignored in the literature. The operators in Eq. (59) generate new interactions not only between the top quark and the physical Higgs field, but also between the top-quark and the neutral(charged) Goldstone boson $G_{0}(G_{\pm})$. The $t\bar{t}G_{0}$ and $tbG_{+}$ vertexes are not modified, but the new vertexes $t\bar{t}G_{0}G_{0}$ and $t\bar{t}G_{+}G_{-}$, and also $t\bar{t}HH$, do appear. In particular, for our calculation, the only quantity that can be affected by these new vertexes is the two-point function $\Sigma(p)$ in Eq. (23), since a new topology with a momentum-independent closed loop of scalars induced by $t\bar{t}HH$ and especially $t\bar{t}G_{0}G_{0}$ and $t\bar{t}G_{+}G_{-}$ can appear. Their contributions are not gauge invariant, also when summed together. This is equivalent to what has been discussed in detail in Ref. Maltoni:2018ttu for the case of modified Higgs-self coupling for the Higgs two-point function (see especially Appendices A and B in that work). However, again similarly to the case discussed in Ref. Maltoni:2018ttu , also for our calculation since the contributions of such diagrams are momentum independent, in the on-shell scheme they do not enter $\delta\psi_{t}$ and are exactly canceled by $\delta m_{t}$, regardless of the momentum of the top-quark. Gauge invariance is therefore preserved.
@@ 6 Higgs boson $H$: Numerical results for scalar one-loop corrections to $t\bar{t}$ distributions
Analogously to Sec. 6, where the case of $S$ was considered, in this section we show and discuss our numerical results for $\sigma_{\rm NP}^{H}$. Also in this case, we do not consider all the higher-order SM corrections introduced in Sec. 2.3 and entering Eqs. (54) and (55), we rather focus on the the relative size of $\sigma_{\rm NP}^{H}$ w.r.t. $\sigma_{\rm LO_{\rm QCD}}$, the LO cross-section associated to the diagrams of Fig. 2. We define the sum of them as $\sigma_{\rm LO_{QCD}+NP}^{H}=\sigma_{\rm LO_{\rm QCD}}+\sigma_{\rm NP}^{H}$.
@@
Considering the others distributions, the behaviour is similar, the differences between the scalar and pseudoscalar corrections are maximal in the area corresponding to the threshold production, namely $p_{T}(t)\simeq 0$, $y(t)\simeq 0$ and $\Delta y(t,\bar{t})\simeq 0$. Also, while the corrections from a purely pseudoscalar $H$ remain always negative, in the case of a purely scalar one they change sign moving towards the tail of $m(t\bar{t})$ and $p_{T}(t)$ distributions. This is also at the origin of the different behaviour of the scalar for large $p_{T}(t)$ w.r.t. large $m(t\bar{t})$. In $t\bar{t}$ production, with no cuts on the phase space, large $m(t\bar{t})$ is dominated by $p_{T}(t)\ll m(t\bar{t})/2$ due to the $t/u$-channel diagrams in the $gg\to t\bar{t}$ process, on the contrary large $p_{T}(t)$ is correlated with $m(t\bar{t})\simeq 2p_{T}(t)$. Thus, in single-differential distributions in $m(t\bar{t})$, for large $m(t\bar{t})$ values there are cancellations between negative and positive contributions from different $p_{T}(t)$ regions, while this dynamics is not present at large values of $p_{T}(t)$, and therefore corrections are larger in absolute value. These features are manifest in Fig. 12, where we show at doubly differential level in $m(t\bar{t})$ and $p_{T}(t)$ the value of the ratio over $\sigma_{\rm NP}/\sigma_{\rm LO_{\rm QCD}}$. The left plot refers to the purely scalar case, while the latter to the purely pseudoscalar one.
@@ 7.1 SM-like pseudodata
In Tab. 1 we report the result for the one-dimensional parameter fit, i.e. separately fitting $\kappa_{t}$ and $\tilde{\kappa}_{t}$. We perform the fit by using for theory predictions both $\sigma^{{\rm mult.},H}_{\rm SM+NP}$ and $\sigma^{{\rm add.},H}_{\rm SM+NP}$ (see Eqs. (55) and Eqs. (55)) quantities. For pseudo data we consistently use in the two cases the prediction for $(\kappa_{t},\tilde{\kappa}_{t})=(1,0)$. Thus, by construction, the fit on the scalar coupling must result in $\kappa_{t}=1$ and indeed we find that. In Tab. 1 we report besides central values the $1\sigma$, $2\sigma$ and $3\sigma$ errors. In fact, by looking at the expression of $\sigma_{\rm NP}^{H}$ in Eq. (51), it is cleat that not only $(\kappa_{t},\tilde{\kappa}_{t})=(1,0)$ but also $(\kappa_{t},\tilde{\kappa}_{t})=(-1,0)$ is a solution of the $\chi^{2}$ function minimisation, and it will be manifest when discussing the plots for the simultaneous $\kappa_{t}$ and $\tilde{\kappa}_{t}$ fits. When the errors become larger than 1, the error intervals for the $(\kappa_{t},\tilde{\kappa}_{t})=(1,0)$ and $(\kappa_{t},\tilde{\kappa}_{t})=(-1,0)$ solutions overlap. This is the origin of the exact $-1$ value in some of the results displayed for the errors. It can also be noted that the $1\sigma$ error in the case of the $\kappa_{t}$ fit is compatible with the ones of Refs. CMS:2019art ; CMS:2020djy . This is precisely due to the strategy that we have described in Sec. 4.1 and that we have adopted in order to obtain a simplified framework for mimicking the accuracy already achieved at the experimental level. The difference between results obtained between the fit in the multiplicative and additive approaches can be regarded as theory uncertainties in modelling predictions for both the SM and NP contributions.
@@ Appendix B Renormalised $t\bar{t}g$ vertex and one-loop $qq\to t\bar{t}$ amplitude
$\bar{u}_{i}(p_{2})\widehat{\Gamma}_{ij}^{\mu,a}v_{j}(p_{1})=(-ig_{s})t^{a}_{ij}\leavevmode\nobreak\ \bar{u}_{i}(p_{2})\left[\widehat{F}_{1}(q^{2})\gamma^{\mu}+i\frac{\sigma^{\mu\nu}}{2m_{t}}p_{\nu}\left(F_{2}(q^{2})+iF_{3}(q^{2})\gamma_{5}\right)\right]v_{j}(p_{1}),$ (74)

### v2

@@ 5 Higgs boson $H$: Theoretical framework
In Sec. 5.1 we provide the relevant Lagrangian, the notation used and the main formulas that are analogous to the ones provided in Sec. 2 for the case of the scalar $S$. In Sec. 5.2 we explain in detail how we have derived such expressions, recycling the results of Sec. 2 for now the Higgs case, highlighting important differences w.r.t. the scalar $S$ case. In Sec. 5.3 we show how this calculation can be reinterpreted in the SMEFT framework, which further support the consistency of our approach.
@@ 5.1 Lagrangian, notation and relevant formulas
$\mathcal{L}_{H,\,{\rm NP}}\equiv-\overline{\psi}_{t}\left[\frac{\left(y_{t}-y_{t}^{\rm SM}\right)}{\sqrt{2}}+i\frac{\tilde{y}_{t}}{\sqrt{2}}\gamma_{5}\right]\psi_{t}H\,\equiv-\frac{y_{t}^{\rm SM}}{\sqrt{2}}\leavevmode\nobreak\ \overline{\psi}_{t}\left[\left(\kappa_{t}-1\right)+i\tilde{\kappa}_{t}\gamma_{5}\right]\psi_{t}H\,,$ (47)
@@
Unlike the case of the scalar $S$, where the contribution from the $t\bar{t}S$ has been taken into account, we will not include the contribution from the $t\bar{t}H$ final state in our studies for the sensitivity on $\kappa_{t}$ and $\tilde{\kappa}_{t}$ in Sec. 7 as that corresponds to a different analysis at the LHC. Moreover, the inclusion of the $t\bar{t}H$ contribution is not necessary for avoiding IR sensitivity, since $m_{H}$ is large. For the same reason, as manifest in Figs. 6 and 8, even if we included such contribution, its impact on our results would be totally negligible.
@@ 5.2 Similarities and differences with the scalar $S$ case
From the pure calculation side the two scenarios are therefore equivalent and via Eqs. (56)–(58) it is possible to obtain $\sigma_{\rm NP}^{H}$ in Eq. (51) from $\sigma_{\rm NP}$ in Eq. (31). However, this cannot be achieved by simply applying them directly as we will explain in the next paragraph. As already said, $\sigma_{\rm NP}^{H}$ corresponds to the renormalised one-loop virtual corrections induced by the diagrams of Fig. 1, with $H$ in the place of $S$, to $\sigma_{\rm LO_{\rm QCD}}$, is the LO cross section from purely QCD interactions (the diagrams in Fig. 2). Renormalisation is again understood in the on-shell scheme, as in the scalar S case.
@@ 5.3 The SMEFT perspective
The consistency of the kappa-framework for such calculation and especially the possibility of recycling the calculation for $S$ of Sec. 2.2 also for the Higgs itself and including not only terms proportional to $(\kappa_{t}-1)^{2}\simeq\Re^{2}(C^{u\Phi}_{tt})$ but also to $\kappa_{t}-1\simeq\Re(C^{u\Phi}_{tt})$ relies on some subtleties that are typically ignored in the literature. The operators in Eq. (59) generate new interactions not only between the top quark and the physical Higgs field, but also between the top-quark and the neutral(charged) Goldstone boson $G_{0}(G_{\pm})$. The $t\bar{t}G_{0}$ and $tbG_{+}$ vertexes are not modified, but the new vertexes $t\bar{t}G_{0}G_{0}$ and $t\bar{t}G_{+}G_{-}$, and also $t\bar{t}HH$, do appear. In particular, for our calculation, the only quantity that can be affected by these new vertexes is the two-point function $\Sigma(p)$ in Eq. (23), since a new topology with a momentum-independent closed loop of scalars induced by $t\bar{t}HH$ and especially $t\bar{t}G_{0}G_{0}$ and $t\bar{t}G_{+}G_{-}$ can appear. Their contributions are not gauge invariant, also when summed together. This is equivalent to what has been discussed in detail in Ref. Maltoni:2018ttu for the case of modified Higgs-self coupling for the Higgs two-point function (see especially Appendices A and B in that work). However, again similarly to the case discussed in Ref. Maltoni:2018ttu , also for our calculation since the contributions of such diagrams are momentum independent, in the on-shell scheme they do not enter $\delta\psi_{t}$ and are exactly canceled by $\delta m_{t}$, regardless of the momentum of the top-quark. Gauge invariance is therefore preserved.
@@ 6 Higgs boson $H$: Numerical results for scalar one-loop corrections to $t\bar{t}$ distributions
Analogously to Sec. 6, where the case of $S$ was considered, in this section we show and discuss our numerical results for $\sigma_{\rm NP}^{H}$. Also in this case, we do not consider all the higher-order SM corrections introduced in Sec. 2.3 and entering Eqs. (54) and (55), we rather focus on the relative size of $\sigma_{\rm NP}^{H}$ w.r.t. $\sigma_{\rm LO_{\rm QCD}}$, the LO cross-section associated to the diagrams of Fig. 2. We define the sum of them as $\sigma_{\rm LO_{QCD}+NP}^{H}=\sigma_{\rm LO_{\rm QCD}}+\sigma_{\rm NP}^{H}$.
@@
Considering the other distributions, the behaviour is similar, the differences between the scalar and pseudoscalar corrections are maximal in the area corresponding to the threshold production, namely $p_{T}(t)\simeq 0$, $y(t)\simeq 0$ and $\Delta y(t,\bar{t})\simeq 0$. Also, while the corrections from a purely pseudoscalar $H$ remain always negative, in the case of a purely scalar one they change sign moving towards the tail of $m(t\bar{t})$ and $p_{T}(t)$ distributions. This is also at the origin of the different behaviour of the scalar for large $p_{T}(t)$ w.r.t. large $m(t\bar{t})$. In $t\bar{t}$ production, with no cuts on the phase space, large $m(t\bar{t})$ is dominated by $p_{T}(t)\ll m(t\bar{t})/2$ due to the $t/u$-channel diagrams in the $gg\to t\bar{t}$ process, on the contrary large $p_{T}(t)$ is correlated with $m(t\bar{t})\simeq 2p_{T}(t)$. Thus, in single-differential distributions in $m(t\bar{t})$, for large $m(t\bar{t})$ values there are cancellations between negative and positive contributions from different $p_{T}(t)$ regions, while this dynamics is not present at large values of $p_{T}(t)$, and therefore corrections are larger in absolute value. These features are manifest in Fig. 12, where we show at doubly differential level in $m(t\bar{t})$ and $p_{T}(t)$ the value of the ratio over $\sigma_{\rm NP}/\sigma_{\rm LO_{\rm QCD}}$. The left plot refers to the purely scalar case, while the latter to the purely pseudoscalar one.
@@ 7.1 SM-like pseudodata
In Tab. 1 we report the result for the one-dimensional parameter fit, i.e. separately fitting $\kappa_{t}$ and $\tilde{\kappa}_{t}$. We perform the fit by using for theory predictions both $\sigma^{{\rm mult.},H}_{\rm SM+NP}$ and $\sigma^{{\rm add.},H}_{\rm SM+NP}$ (see Eqs. (55) and Eqs. (55)) quantities. For pseudo data we consistently use in the two cases the prediction for $(\kappa_{t},\tilde{\kappa}_{t})=(1,0)$. Thus, by construction, the fit on the scalar coupling must result in $\kappa_{t}=1$ and indeed we find that. In Tab. 1 we report besides central values the $1\sigma$, $2\sigma$ and $3\sigma$ errors. In fact, by looking at the expression of $\sigma_{\rm NP}^{H}$ in Eq. (51), it is clear that not only $(\kappa_{t},\tilde{\kappa}_{t})=(1,0)$ but also $(\kappa_{t},\tilde{\kappa}_{t})=(-1,0)$ is a solution of the $\chi^{2}$ function minimisation, and it will be manifest when discussing the plots for the simultaneous $\kappa_{t}$ and $\tilde{\kappa}_{t}$ fits. When the errors become larger than 1, the error intervals for the $(\kappa_{t},\tilde{\kappa}_{t})=(1,0)$ and $(\kappa_{t},\tilde{\kappa}_{t})=(-1,0)$ solutions overlap. This is the origin of the exact $-1$ value in some of the results displayed for the errors. It can also be noted that the $1\sigma$ error in the case of the $\kappa_{t}$ fit is compatible with the ones of Refs. CMS:2019art ; CMS:2020djy . This is precisely due to the strategy that we have described in Sec. 4.1 and that we have adopted in order to obtain a simplified framework for mimicking the accuracy already achieved at the experimental level. The difference between results obtained between the fit in the multiplicative and additive approaches can be regarded as theory uncertainties in modelling predictions for both the SM and NP contributions.
@@ Appendix B Renormalised $t\bar{t}g$ vertex and one-loop $qq\to t\bar{t}$ amplitude
$\bar{u}_{i}(p_{2})\widehat{\Gamma}_{ij}^{\mu,a}v_{j}(p_{1})=(-ig_{s})t^{a}_{ij}\leavevmode\nobreak\ \bar{u}_{i}(p_{2})\left[\widehat{F}_{1}(q^{2})\gamma^{\mu}+i\frac{\sigma^{\mu\nu}}{2m_{t}}q_{\nu}\left(F_{2}(q^{2})+iF_{3}(q^{2})\gamma_{5}\right)\right]v_{j}(p_{1}),$ (74)